If A is the midpoint of line MT and T is the midpoint of line MH then MA =TH
A) True point T is between MH and seperates the segment into 2 equal length
B) False there is no shared midpoint between line MA and TH

Answer by richwmiller(17219)

MA is 1/2*(MT) and MT=TH therefore MA=1/2*TH not MA=TH
Therefore B) False